OpenStudy (anonymous):

!!!!!!!!!!!MEDAL AND FAN Jane has 38 marbles of the same shape and size in a box. The colors and amounts of the marbles are shown below: 14 white marbles 9 gray marbles 15 blue marbles Without looking in the box, Jane takes out a marble at random. She then replaces the marble and takes out another marble from the box. What is the probability that Jane takes out a gray marble in both draws? 9/38 * 8/37= 72 /1406 9 / 38 + 8/ 37 = 637 /1406 9/38 + 9/38 = 684 /1444 9/38 * 9 /38 =81/1444

OpenStudy (anonymous):

i think 3 but idk

OpenStudy (anonymous):

@Michele_Laino ??

OpenStudy (jango_in_dtown):

its d

OpenStudy (jango_in_dtown):

both the events are independent so P(AB)=P(A)*P(B)=(9/38)*(9/38)

OpenStudy (anonymous):

Thank you so much !!